During a baseball game a batter hits a long flyball. If the ball remains in the air for 6.0 sec, how high does it rise?
1
Expert's answer
2012-01-31T11:00:47-0500
When the ball is at its high point, it has zero velocity for an instant and falls for half the time it was in the air, the distance it falls is:

H = 0.5g(T/2)² = 0.5(9.8m/s²)(6/2s)² = 44.1 m.